A little help please!?
Snack mix is made with peanuts and M&M's and must make at least 100 pounds to be packaged and sold. The mixture must contain at least 40 pounds of peanuts and must contain at least 20 pounds of M&M's. Peanuts cost $4 per pound and M&M's cost $3 per pound. How many pounds of peanuts and how many pounds of M&M's would minimize the total cost of the snack mix?

Thank you!

Answer by stanbon(75887)

Snack mix is made with peanuts and M&M's and must make at least 100 pounds to be packaged and sold.
--------------------
The mixture must contain at least 40 pounds of peanuts and must contain at least 20 pounds of M&M's.
--------------------
Peanuts cost $4 per pound and M&M's cost $3 per pound.
-------------------------
How many pounds of peanuts and how many pounds of M&M's would minimize the total cost of the snack mix?
---------
m >= 20
p >= 40
m+p >=100
Cost = 3m + 4p
----
with (20,80) cost = 3*20 + 4*80 = 60 + 320 = 380 dollars
with (60,40) cost = 3*60 + 4*40 = 180 + 160 = 340 dollars
----
Minimum cost comes with 60 lbs of M&M and 40 lbs of peanuts.
================================================================
Cheers,
Stan H.
============